Two investments totaling $15,500 produce an annual income of $1450. One investment yields 10% per year, while the other yields 5% per year. How much is invested at each rate?
Answer by VFBundy(438)

10% rate:
Principal = p
Rate = 0.10
Interest = 0.10p

5% rate:
Principal = 15500 - p
Rate = 0.05
Interest = 0.05(15500 - p) = 775 - 0.05p

0.10p + (775 - 0.05p) = 1450

0.05p + 775 = 1450

0.05p = 675

p = 13500

10% rate:
Principal = p = $13,500

5% rate:
Principal = 15500 - p = 15500 - 13500 = $2000